(3x)^2/3 = 16
I know the answer is x=64/3
Can someone please explain how they got the answer.
Thank you,
Trey

Answer by Alan3354(69443)

(3x)^2/3 = 16
-------
Raise both sides to the 3rd power
(3x)^2 = 4096
Take the sqrt
3x = 64
x = 64/3